OpenStudy (swaqqout_kid1):

As inflation increases, the value of the dollar decreases at the same rate. The value of the dollar after t months of inflation can be found by raising the difference of 1 and the rate of inflation, r, to the power of t. If the current rate of inflation is 4%, what will the value of the dollar, d, be after 8 months? Round to the nearest cent. @sangya21

OpenStudy (anonymous):

d = (1-r)^t d =(1-4%)^8